A psychology professor claims that the variance of IQ scores for college students is σ^2 = 100. Let X1, X2,…, X23 be a random sample of n = 23 IQ scores and assume these scores come from a N(µ, σ^2) distribution. Let S^2 = (1/22) ∑_(i=1 to 23) (Xi – X bar)^2 be the sample variance of these scores. The observed value of S^2 = 147.82. Show that the Var(S^2) = 10000/11. Thus, the standard deviation of S^2 is 30.15.
Answer #1

By defa of x, xn tor(0,0%), (0-1) 5² ~ x² na SE [name] =na, var [ S] =2camy ne var [ 5²7= 204 (0-1) stere n=23,6²2100 so thatproperties of chi square distribution is used. For further query in above, comment.